A machine has four components, A, B, C, and D, set up in such a manner that all four parts must work for the machine to work properly. Assume the probability of one part working does not depend on the functionality of any of the other parts. Also assume that the probabilities of the individual parts working are P(A) = P(B) = 0.93, P(C) = 0.95, and P(D) = 0.92. Find the probability that the machine works properly.

A)

0.8128

B)

0.2441

C)

0.8217

D)

0.7559
I dont even know what route to take or where to start on this one..please help at least point me in the right direction!

Answer by alicealc(293)

because the all of the four parts must work, the probability that the machine works = 0.93 * 0.93 * 0.95 * 0.92 = 0.7559 (D)